A 5.60-kg particle moves along the x axis. Its position varies with time according to x = t + 4.0t3, where x is in meters and t is in seconds.

(a) Find the kinetic energy of the particle at any time t. (Use the following as necessary: t.)
K =  

(b) Find the magnitude of the acceleration of the particle and the force acting on it at time t. (Use the following as necessary: t.)
a =
F =

(c) Find the power being delivered to the particle at time t. (Use the following as necessary: t.)
P =  

(d) Find the work done on the particle in the interval t = 0 to t = 4.30 s.
J
Answer #1

x = t + 4.0t^3
V=differentiation(x)=1+12*t^2
a=24t
a)Kinetic energy=0.5*m*v^2
          =0.5*5.6*(1+12*t^2)^2
          =2.8*(1+12*t^2)^2
b)a=24t
F=Ma=5.6*24t=134.4t
c)power=FV=(134.4t)(1+12*t^2)
          134.4t+1612.6*t^3
d)work done=

media%2F573%2F57351381-9b6e-4676-90c4-20